OpenStudy (anonymous):

Shannon is managing a small zoo and has been analyzing the attendance data. Shannon has found that the number of visitors increases exponentially as the temperature increases. Shannon has also found a linear equation that models the number of people who leave the park early depending on the temperature. Describe how Shannon can combine these two functions into a new function and explain what that function would predict.

OpenStudy (anonymous):

Please help? I dont know how to do this without numbers

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Agl202 Thanks so much

OpenStudy (agl202):

Give me a sec

OpenStudy (anonymous):

Okay, thanks :)

OpenStudy (agl202):

Well we can think up something. Above 80° 1 animal asleep Above 85° 2 animals asleep Above 90° 4 animals asleep Above 95° 8 animals asleep Hmm there's a pattern developing. sleeping animals = 2^ (Air Temp - Temp 80° F) Air temp 80° - then 2^(80-80) = 2^0 = 1

OpenStudy (anonymous):

Hmm, so what does this mean exactly? Im confused. D;

OpenStudy (anonymous):

@Agl202 can you help me please

OpenStudy (agl202):

wait @SweetKiss

OpenStudy (anonymous):

I see the pattern is increased bt 5 degrees o:

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

sorry

OpenStudy (agl202):

People leaving the park early = 2 * ((Air Temp -75°) / 5)

OpenStudy (agl202):

Do u think u can do the rest? the remaining thing to do is : "Describe how Shannon can combine these two functions into a new function and explain what that function would predict."

OpenStudy (anonymous):

Okay, so the sleeping animals is equivalent to the people? I'm a tad confused. like what does the sleeping animals have to with it? lol How do I use that as an example? Sorry Im a bit slow :/

OpenStudy (anonymous):

And Im not sure what it means by combined them

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

What function am I combining? There is no numbers, and how would I combine them? is oene of the functions the 2 * (-75/5)?

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

@osamabinswaggin

OpenStudy (anonymous):

@ahamed.